# CMT: Mid-Training for Efficient Learning of Consistency, Mean Flow, and Flow Map Models

### v1

@@ B.1 CIFAR-10, AFHQv2, and FFHQ
B.2 ImageNet 64×64
B.2.1 Experimental Details
@@ B.3 ImageNet 256×256
B.4 ImageNet 512×512
B.4.1 Experimental Details
@@ B.4.2 Cost Details
Appendix C More Experimental Results
C.1 CIFAR-10: Importance of Mid-Training
We further validate the importance of our proposed mid-training using the CIFAR-10 dataset. Previously, our post-training ECT had a linearly decaying learning rate and fixed $\Delta t$, which is different from the original ECT. Therefore, we make the post-training ECT and the vanilla ECT share the same hyperparameter to test the effect of mid-training more fairly. We make the following changes to the setting while keeping other hyperparameters not mentioned unchanged.
  • Vanilla ECT: We set ECT with constant $\Delta t=1/256$ under the 51.2Mimgs budget. The final 1-step / 2-step FID is 3.54 / 2.12.
  • CMT1: Short mid-training for 1.28Mimgs + Long post-training for 49.92Mimgs. Constant $\Delta t=1/256$. The final 1-step / 2-step FID is 3.42 / 2.11.
  • CMT2: Long mid-training for 25.6Mimgs + Short post-training for 25.6Mimgs. Constant $\Delta t=1/256$. The final 1-step / 2-step FID is 3.30 / 2.04.
All models are optimized with a constant 1e-4 learning rate. The results demonstrated that longer mid-training outperforms shorter one, which is in turn better than no mid-training. This further validates the importance of mid-training.
C.2 Evaluating Alternatives for Post-Training Initialization
Ablation Study on Knowledge Distillation (KD) and Slow CMT. Even though we have demonstrated the efficiency of CMT as a mid-training method, other variants are possible. We take the flow map learning ${\bm{\Psi}}_{t\to 0}$ as representative and highlight two straightforward alternatives. Following the design principles of the mid-training stage, these variants are expected to be stable and easy to train (e.g., using stop-gradient-free regression targets), while relying on fewer ad-hoc tricks and hyperparameters:
$\mathcal{L}_{\mathrm{var}}^{(1)}({\bm{\theta}}):=\mathbb{E}_{t,\,p_{t}}\!\left[\;\|{\mathbf{f}}_{\bm{\theta}}({\mathbf{x}}_{t},t)-\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})\|_{2}^{2}\;\right],\quad\mathcal{L}_{\mathrm{var}}^{(2)}({\bm{\theta}}):=\mathbb{E}_{p_{\mathrm{prior}}}\!\left[\;\|{\mathbf{f}}_{\bm{\theta}}({\mathbf{x}}_{T},T)-\hat{\mathbf{x}}_{0}({\mathbf{x}}_{T})\|_{2}^{2}\;\right],$
where $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ denotes the estimate at t=0t{=}0 obtained by running the solver with the pre-trained diffusion model starting from the forward-perturbed sample ${\mathbf{x}}_{t}$ at time $t$, and $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{T})$ is obtained similarly by running backward from a prior sample ${\mathbf{x}}_{T}$, as in Luhman & Luhman (2021).
We call $\mathcal{L}_{\mathrm{var}}^{(1)}({\bm{\theta}})$ Slow CMT as it is doing CMT in principle but omits using intermediate points of the ODE-solver generated trajectory. In contrast, this Slow CMT $\mathcal{L}_{\mathrm{var}}^{(1)}({\bm{\theta}})$ only uses the ODE trajectory end point. Thus, to produce the same amount of data, Slow CMT require more costly ODE-solver inference. Meanwhile, the loss function $\mathcal{L}_{\mathrm{var}}^{(1)}({\bm{\theta}})$ is called Knowledge Distillation (KD) following Luhman & Luhman (2021).
We compare CMT’s mid-training with KD and Slow CMT. KD explicitly focuses on the mapping from noise to the corresponding data only, while CMT focuses on the entire ODE trajectory consistency. We use the CIFAR-10 dataset and train KD and Slow CMT with the same settings as CMT. Subsequently, we use ECT to post-train for flow map learning with the three different initializations for 12.8 Mimgs cost.
Comparing CMT with KD, CMT’s 1/2 step FID (2.74/1.97) is much better than KD’s (3.54/2.19), verifying the benefits of learning intermediate steps.
Comparing CMT with its slow variant, they achieve a similar FID, with CMT’s 1/2 step FID (2.74/1.97) and Slow CMT’s 1/2 step FID (2.75/1.98). However, the Slow CMT’s mid-training stage costs 3x more GPU time since generating the regression target in the Slow version is more costly due to the failure to use intermediate points by the ODE-solver.
C.3 Samples Generated by CMT on ImageNet $\times512$
Appendix D Training Speed and Memory Cost
D.1 Empirical Runtime Comparison
@@
D.2 Analysis of Computational Cost in CM
@@
In short, we analyze the cost per input–target pair. In CMT, one teacher-generated trajectory yields many pairs by matching intermediate states to the clean target, whereas CD and CT generate each pair independently from a sampled time. Per pair, CMT needs one teacher call plus one student forward and one backpropagation; CD needs one teacher call plus the same student cost; CT needs no teacher call but two student forwards and one backpropagation. Hence, CMT is roughly as costly as CD and only slightly slower than CT, consistent with our empirical findings (Appendix D.1). Meanwhile, CMT achieves near–unit teacher cost per pair while using a single student forward, making it a lightweight and effective choice for the mid-training stage.
Appendix E Discussion on CMT Loss Metric

### v2

@@ B.2 CIFAR-10, AFHQv2, and FFHQ
B.3 ImageNet 64×64
B.3.1 Experimental Details
@@ B.4 ImageNet $\times256$
B.5 ImageNet $\times512$
B.5.1 Experimental Details
@@ B.5.2 Cost Details
B.6 MS-COCO Text-to-Image (T2I)
The SiT teacher model is trained with REPA (Yu et al., 2025) for 50k iterations, achieving a 50-step FID of 4.73 using a second-order Heun ODE solver, and the CFG scale is 2. All other models are also trained with REPA. The SiT teacher additionally serves as the pre-training stage.
Regarding MF from scratch, we mainly reuse the default setting in our ImageNet $\times256$ MF experiment. But we set the MF’s effective CFG scale to 2 and the CFG trigger interval as $[0,1]$ since the SiT teacher uses the same CFG hyperparameters. We specifically set $\kappa=0.5$ and $\omega=1$. Regarding MF initialized by SiT and CMT with MF, we use the same MF training configuration for post-training.
For mid-training in CMT, we generate the reference ODE trajectory with the pre-trained SiT with an 8-step Heun solver. We also use a constant learning rate of 1e-4 and do not use any weighting trick, and use squared $\ell_{2}$ loss. We use four random samples to generate trajectories, and each sample provides 28 pairs of $(\hat{{\mathbf{x}}}_{t_{i}},\hat{{\mathbf{x}}}_{t_{j})}$, inducing a total $28\times 4=112$ batch size per iteration.
Appendix C Training Speed and Memory Cost
C.1 Empirical Runtime Comparison
@@
C.2 Analysis of Computational Cost in CM
@@
In short, we analyze the cost per input–target pair. In CMT, one teacher-generated trajectory yields many pairs by matching intermediate states to the clean target, whereas CD and CT generate each pair independently from a sampled time. Per pair, CMT needs one teacher call plus one student forward and one backpropagation; CD needs one teacher call plus the same student cost; CT needs no teacher call but two student forwards and one backpropagation. Hence, CMT is roughly as costly as CD and only slightly slower than CT, consistent with our empirical findings (Appendix C.1). Meanwhile, CMT achieves near–unit teacher cost per pair while using a single student forward, making it a lightweight and effective choice for the mid-training stage.
Theoretical NFEs Comparison with CMT and Slow CMT Variant.
Mathematically, both the CMT loss $\mathcal{L}_{\text{CMT-CM}}$ in Equation 7 and the Slow CMT variant $\mathcal{L}_{\mathrm{var}}^{(1)}$ are discrete-time approximations to the same oracle loss $\mathcal{L}_{\text{oracle-CM}}$ in Equation 5. However, they differ crucially in how the teacher-generated supervision targets are constructed. In CMT, we first draw $\mathbf{x}_{T}\sim p_{\mathrm{prior}}$, run the teacher ODE solver once from $T$ down to 0 to obtain a full trajectory $\{\hat{\mathbf{x}}_{t_{i}}\}_{i=0}^{M}$, and then use the single endpoint $\hat{\mathbf{x}}_{t_{0}}$ as a shared target for all intermediate states $(\hat{\mathbf{x}}_{t_{i}},t_{i})$ on that trajectory. Thus, one teacher solve yields $M$ supervised pairs $(\hat{\mathbf{x}}_{t_{i}},t_{i})\mapsto\hat{\mathbf{x}}_{t_{0}}$. In contrast, in Slow CMT (based on $\mathcal{L}_{\mathrm{var}}^{(1)}$), we sample $(\mathbf{x}_{t},t)$ and, for each such pair, independently run the teacher solver from $\mathbf{x}_{t}$ to $t=0$ to obtain $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$. Each supervised pair $(\mathbf{x}_{t},t)\mapsto\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ therefore requires its own teacher integration. The two objectives coincide in the continuous-time limit, but their discrete implementations use teacher trajectories in fundamentally different ways.
Because of this difference, the mid-training cost of Slow CMT is significantly higher for a fixed number of supervised pairs. Let $M$ denote the number of discretization steps on $[0,T]$, and let $\mathrm{NFEs}_{\text{traj}}$ be the teacher function evaluations required for one trajectory. In CMT, a single teacher trajectory is amortized over $M$ pairs, so the teacher cost per training pair is $\mathrm{NFEs}_{\text{traj}}/M=\mathcal{O}(1)$. In Slow CMT, each training pair requires a fresh teacher solve, so the teacher cost per pair is $\mathcal{O}(\mathrm{NFEs}_{\text{traj}})$. Consequently, to train on the same number of supervised pairs, Slow CMT would require roughly a factor of $M$ more teacher evaluations (and thus GPU time) than CMT; conversely, under a fixed teacher-compute budget, Slow CMT can only afford about $1/M$ as many pairs. In our experimental configuration, this leads to an overhead of roughly $3\times$ in wall-clock GPU time, which is what we refer to in the paper when stating that the Slow CMT mid-training stage is about three times more expensive. We did not perform an exhaustive sweep that precisely equalizes GPU hours between the two variants during the review period, but the existing experiments already show that CMT achieves better FID at substantially lower teacher cost, which is exactly the mid-training efficiency advantage we aim to demonstrate.
Appendix D Discussion on CMT Loss Metric